OpenStudy (anonymous):

Write the equation of the circle with center (-6, -4) and containing the point (-2, -1).

OpenStudy (anonymous):

\[ (x+6)^2 +(y+4)^2 =r^2\\ (-2+6)^2 +(-1+4)^2 = r^2=25\\ (x+6)^2 +(y+4)^2 =25 \]